# Branching Flows: Discrete, Continuous, and Manifold Flow Matching with Splits and Deletions

### v1

@@ 3.3 Sampling from the Conditional Path
For training, we need samples at time $t$ from the conditional path $\tilde{X}_{t}\sim p_{t\mid Z}$. With Branching Flows, even though the element evolution along a branch for the conditional paths is independent of the other elements, because of the branching structure the distribution of the element value at any given time is not. This prevents us from adopting the usual strategy for flow matching which draws conditional samples for each element independently.
However we can efficiently sample the waiting time until the next split or deletion event. If we additionally have an efficient per-element sampler for $p(\tilde{x}_{v}^{(k)}\mid\tilde{x}_{s}^{(k)},a_{k})$, where $s<v$ and $a_{k}$ is the anchor associated with the node ahead of $\tilde{x}_{s}^{(k)}$, then we can sample from the conditional path at $t$ by recursively sampling the waiting time until the next split or deletion event, and then sampling the value of the element at the next split event (or at $t$ if the next split event is after $t$).
For the base process in our empirical examples below, we use an Ornstein–Uhlenbeck (OU) like process with time-inhomogeneous diffusion, and with mean-reversion towards the anchors, where the endpoint conditioned distribution $p(\tilde{x}_{v}^{(k)}\mid\tilde{x}_{s}^{(k)},a_{k})$ follows from the results in AlbanoGiorno2020OU. For our discrete states we use an instance of Discrete Flow Matching (DFM) corresponding to equation 10 in Gat2024DiscreteFM, which includes unmasking and uniform noise.
Figure 2: Branching Flows construction. Left outlines the sampling of $Z$ when $x_{1}$ has 4 elements and $x_{0}$ has two elements, for a conditional bridge that will incur a single deletion. Right depicts the conditional sampling of $x_{t}|Z$, where elements can split and be deleted but only, in the conditional path, according to the pre-determined branching structure of the two trees $\mathcal{T}$ in $Z$.
@@ 3.4 Parameterization of the Generator and Loss
The base element generator is constructed as desired. For the split rate and the deletion rate, just as described for Counting Flows and Deletion Flows above, we frame these as $x_{1}$-prediction, and have the model predict the expected number of splits ahead of the current element by $t=1$, and the probability of deletion by $t=1$.
@@
The Conditional Branching Flows (CBF) loss is the sum of three Bregman divergences: one for the base process, one against the split intensity, and one against the deletion intensity:
$L_{\mathrm{CBF}}(\theta)=\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},\tilde{X}_{t}\sim p_{t|Z}(d\tilde{x}|z)}\Big[$ $D^{\mathrm{split}}_{t,\tilde{X}_{t}}(R_{t}^{Z}(\tilde{X}_{t}),R_{t}^{\theta}(\tilde{X}_{t}))$
$+D_{t,\tilde{X}_{t}}^{\mathrm{del}}(\rho_{t}^{Z}(\tilde{X}_{t}),\rho_{t}^{\theta}(\tilde{X}_{t}))$
$+D_{t,\tilde{X}_{t}}^{\mathrm{base}}\big(F_{t}^{Z,\mathrm{base}}(\tilde{X}_{t}),F_{t}^{\theta,\mathrm{base}}(\tilde{X}_{t})\big)\Big].$
[…]
The loss for the splitting increments mirrors that from Counting Flows (section 2.3.1), and deletions from Deletion Flows (section 2.3.2), while the base process depends on the specific application. Our chosen loss is the following:
$L_{\mathrm{CBF}}(\theta)=\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},\tilde{X}_{t}\sim p_{t|Z}(d\tilde{x}|z)}\Big[$ $\sum_{i=1}^{N}\Big({R_{t,i}^{\theta}(\tilde{X}_{t})-R_{t,i}^{Z}(\tilde{X}_{t})\log R_{t,i}^{\theta}(\tilde{X}_{t})}\Big)$ (1)
$+\sum_{i=1}^{N}\Big(-[\rho^{Z}_{t,i}(\tilde{X}_{t})\log\rho_{t,i}^{\theta}(\tilde{X}_{t})+(1-\rho^{Z}_{t,i}(\tilde{X}_{t}))\log(1-\rho_{t,i}^{\theta}(\tilde{X}_{t}))]\Big)$
$+\sum_{i=1}^{N}\Big(D_{t,i}^{\mathrm{base}}\big(F_{t,i}^{a(i),\mathrm{base}}(\tilde{X}_{t}),F_{t,i}^{\theta,\mathrm{base}}(\tilde{X}_{t}\big)\Big)\Big].$
Here, $a(i)$ denotes the anchor associated with the branch specified by the branch indicator of the $i$-th component of the augmented state $\tilde{X}_{t}$.
3.5 Training
@@
Training proceeds exactly as is standard in flow matching, by stochastic gradient descent. We draw $t\sim\mathcal{D}$, $X_{1}\sim q$, $Z\sim p_{Z\mid X_{1}}$, and then draw $\tilde{X}_{t}\sim p_{t\mid Z}$. From $Z,\tilde{X}_{t}$, using the branch indicator of each element, we recover $R^{Z}_{t,i}(\tilde{X}_{t})$, $\rho^{Z}_{t,i}(\tilde{X}_{t})$, and $F^{a(i),\mathrm{base}}_{t,i}(\tilde{X}_{t})$ and take gradient steps on the loss in equation (1).
3.6 Sampling from the Marginal Path
@@
We sample from the marginal paths via Euler steps. A trained model outputs element-wise predictions $R^{\theta}_{t,i}(\tilde{X}_{t})$, $\rho^{\theta}_{t,i}(\tilde{X}_{t})$, and $F^{\theta,\mathrm{base}}_{t,i}(\tilde{X}_{t})$. For a step size $\Delta_{t}$, we first take a step on the base process according to the model’s expected generator. For the branching and deletion part of each step, we calculate per-element split intensities $h_{\operatorname{split}}(t)\cdot R^{\theta}_{t,i}(\tilde{X}_{t})$ and deletion intensities $h_{\operatorname{del}}(t)\cdot\rho^{\theta}_{t,i}(\tilde{X}_{t})$, and we sample which elements split or delete within the step. Elements with deletion events are removed, and elements with split events are duplicated in-place.
3.7 Modeling Branch Indices
Under Augmented Generator Matching, the conditional process must be Markov in $\tilde{X}_{t}=(X_{t},G_{t})$. Since the conditional generator is conditioned on $G_{t}=g_{t}$, the model parameterization of the generator must depend on $t,x_{t},g_{t}$, which means that it must explicitly consider the branch indices. This seems architecturally complex at first, since these are variable length sequences, one per element of $x_{t}$. But since the sequence is ordered (and the model already knows the sequence order via positional encoding) and the trees are planar, knowing the tree and branch indicators is equivalent to knowing, between any two elements, the number of tree nodes that they have in common along the path back to $t=0$. This is simply a pairwise count matrix, which can be exposed to the model via e.g. pair features, or pair bias terms in the transformer attention mechanism.
We have found, however, that this is not at all necessary, and having the model take $t,x_{t}$ is sufficient, so long as the state is associated with the correct branch during the conditional process (to link it to the correct anchor for the training objective). Dropping explicit branch tracking simplifies the model architecture and the bookkeeping required during inference, which would otherwise have to track the tree that is constructed by the sequence of split and deletion events.
So while under Generator Matching one can construct theory-satisfying Branching Flows by constructing count matrices and passing these to the model, all of our empirical results below use this simplified form where we drop the branch tracking index from the model. A forthcoming BillerazGM more general version of the discrete-only theorems in Havasi2025EditFF clarifies that dropping the tree and branch indicators amounts to a valid marginal generator for $X_{t}$ that marginalizes over the time-dependent latent process $G_{t}$. For the case where $X_{t}$ is discrete (e.g. the antibody sequence generation example below), this is already established by the theorems in Havasi2025EditFF.
3.8 Concrete Schemes for Sampling $Z$

### v3

@@ 3.3 Sampling from the Conditional Path
[…]
For the base process in our empirical examples below, we use an Ornstein–Uhlenbeck (OU) like process with time-inhomogeneous diffusion, and with mean-reversion towards the anchors, and for discrete states we use an instance of Discrete Flow Matching (DFM) corresponding to equation 10 in Gat2024DiscreteFM, which includes unmasking and uniform noise. See app. D for details.
@@ 3.4 Parameterization of the Generator and Loss
The Conditional Branching Flows (CBF) loss (cf. Appendix E.4), which uses the conditional process to train a model that samples from the marginal path, is the sum of three Bregman divergences: one for the base process, one against the split intensity, and one against the deletion intensity:
$L_{\mathrm{CBF}}(\theta)=$ $\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},(X_{t},G_{t})\sim p_{t|Z}(d\tilde{x}|z)}$
$\Big[D^{\mathrm{split}}_{t,X_{t}}(R_{t}^{Z,G_{t}}(X_{t}),R_{t}^{\theta}(X_{t}))$
$+D_{t,X_{t}}^{\mathrm{del}}(\rho_{t}^{Z,G_{t}}(X_{t}),\rho_{t}^{\theta}(X_{t}))$
$+D_{t,X_{t}}^{\mathrm{base}}\big(F_{t}^{Z,G_{t},\mathrm{base}}(X_{t}),F_{t}^{\theta,\mathrm{base}}(X_{t})\big)\Big].$
Concretely, we use an elementwise loss where we parameterize the split and deletion rates in terms of the number of splits by $t=1$ and the probability of deletion by $t=1$ (i.e. $X_{1}$-prediction”), denoting $R_{t,i}^{Z,G_{t}}(X_{t})$ for the number of split increments ahead of the $i$’th component, and $\rho_{t,i}^{Z,G_{t}}(X_{t})\in\{0,1\}$ for the $i$’th component of $\rho_{t}^{Z,G_{t}}(X_{t})$. These are parametrized by $R_{t,i}^{\theta}(X_{t})\in(0,\infty)$ and $\rho_{t,i}^{\theta}(X_{t})\in(0,1)$ respectively. We write $F_{t,i}^{Z,G_{t},\mathrm{base}}(X_{t})$ as the ‘base’ generator of a per-element conditional process along a single branch, steered to terminate (by $t=1$) at the anchor $a(i,Z,G_{t})$. This is the anchor at the end of the branch along which the $i$-th component of $X_{t}$ is evolving (tracked by the branch indicator of the $i$-th component of the augmented state $\tilde{X}_{t}$). Our loss is then:
$L_{\mathrm{CBF}}(\theta)=\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},(X_{t},G_{t})\sim p_{X_{t},G_{t}|Z}(\mathrm{d}x,dg|z)}$ (1)
$\qquad\Big[\sum_{i=1}^{L_{t}}\Big({R_{t,i}^{\theta}(X_{t})-R_{t,i}^{Z,G_{t}}(X_{t})\log R_{t,i}^{\theta}(X_{t})}\Big)$
$\qquad+\sum_{i=1}^{L_{t}}\Big(-[\rho^{Z,G_{t}}_{t,i}(X_{t})\log\rho_{t,i}^{\theta}(X_{t})$
$\qquad+(1-\rho^{Z,G_{t}}_{t,i}(X_{t}))\log(1-\rho_{t,i}^{\theta}(X_{t}))]\Big)$
$\qquad+\sum_{i=1}^{L_{t}}\Big(D_{t,X_{t},i}^{\mathrm{base}}(F_{t,i}^{Z,G_{t},\mathrm{base}}(X_{t}),F_{t,i}^{\theta,\mathrm{base}}(X_{t}))\Big)\Big].$
Note that only $X_{t}$ (and not $G_{t}$) is input to the model, as $G_{t}$, like $Z$, is marginalized out in the training process per app. E.3.
3.5 Training and Sampling
@@
Training proceeds exactly as is standard in flow matching, by stochastic gradient descent. We draw $t\sim\mathcal{D}$, $X_{1}\sim q$, $Z\sim p_{Z|X_{1}}$, and then draw $\tilde{X}_{t}=(X_{t},G_{t})\sim p_{t\mid Z}(d\tilde{x}_{t}|z)$. From $Z,X_{t}$, and $G_{t}$, we recover $R^{Z,G_{t}}_{t,i}(X_{t})$, $\rho^{Z,G_{t}}_{t,i}(X_{t})$, and $F^{Z,G_{t},\mathrm{base}}_{t,i}(X_{t})$, and take gradient steps on the loss in equation (1).
@@
We sample from the marginal paths by taking small steps in time. A trained model outputs element-wise predictions $R^{\theta}_{t,i}(X_{t})$, $\rho^{\theta}_{t,i}(X_{t})$, and $F^{\theta,\mathrm{base}}_{t,i}(X_{t})$. For a step size $\Delta t$, we first take a step on the base process according to the model’s learned expected generator, independently per element, just as in standard Generator Matching. For the branching and deletion part of each step, we calculate per-element split intensities $h_{\operatorname{split}}(t)\cdot R^{\theta}_{t,i}(X_{t})$ and deletion intensities $h_{\operatorname{del}}(t)\cdot\rho^{\theta}_{t,i}(X_{t})$, and we sample which elements split or delete within the step. Elements with deletion events are removed, and elements with split events are duplicated in-place. The trees and branch indicators have been marginalized out and are not considered when sampling.
4 Results
We consider three datasets: i) small molecules, where we jointly generate atom positions and atom labels, ii) Antibody heavy chains (discrete sequence only), and iii) protein backbones, comprising amino acid ‘frames’ (jumper2021highly) with Euclidean positions & SO(3) orientations, and discrete amino acid sequence labels. We are interested in understanding the degree to which Branching Flows-trained models can learn the data distribution, via comparisons of domain-specific property projections from the sample space. We also use UMAP (mcinnes2020umapuniformmanifoldapproximation) (or seqUMAP, hanke2022multivariate for the antibody sequence case) to visualize the sample and data distribution.
4.1 QM9
Figure 3: QM9 Branching Flows sampling trajectories. A visualization of inference trajectories from a QM9-trained Branching Flows model. The final sampled molecule, when $t=1$, is depicted on the right. At every inference step from $t=0$ to $t=1$, we draw the current $x_{t}$ state as colored points, with a small rightward displacement, which shows the branching and deletion history as $x_{t}$ is transported from a single element when $t=0$ to the final sampled molecule. The color of the trails shows the atom type which begins as ‘Mask’ (in orange) and switches to concrete atoms as $t\rightarrow 1$.
Table 1: Distribution agreement ($1-\mathrm{KS}_{D}$, higher is better) between properties of samples from each model vs. QM9 data. From left to right: Molecular weight, LogP, hydrogen-bond donors, hydrogen-bond acceptors, number of: carbons, hydrogens, oxygens, nitrogens, fluorines, and the total number of atoms.
Model mw logp hbd hba C H O N F atoms
Figure 4: QM9 data vs generated samples. 10,000 random samples from the QM9 dataset vs 10,000 samples generated using a Branching Flows model. Panel A computes molecular fingerprints (using RDKit) and jointly embeds QM9 data and generated samples using UMAP. Left shows QM9-only embeddings, colored by the number of carbon atoms in each molecule; right shows generated samples, similarly colored, and center shows the overlap of QM9 (blue) vs Branching Flows generated samples (red), showing similar distributions. Panel B depicts property distributions from QM9 and Branching Flows generated samples, including atom counts, as well as a number of standard molecular descriptors computed with RDKit.
QM9 (ramakrishnan2014quantum) contains small organic molecules selected from the GDB (Generated Database) chemical space enumeration. We train Branching Flows to jointly generate atom coordinates and discrete labels; full data preprocessing, Branching Flows specification, and model/training details are in Appendix F.1.
Figure 3 shows typical inference trajectories for QM9 molecules. Starting from a single $x_{0}$ element with a random position and a masked label, the process undergoes a sequence of bifurcations and deletions governed by the rates output from the model, with the discrete states switching along the trajectory as well, terminating at a generated molecule. The only relevant benchmark for our purposes is whether the model learns, via Branching Flows, to sample from the data distribution. To investigate this, we generated 10k samples and compared these to 10k QM9 samples (See appendix F.1 for details). Figure 4 compares generated and data distributions for molecule fingerprints, which encode molecule properties into a fixed-length bit vector, embedded into two-dimensions by UMAP, showing no apparent differences. Also shown are histograms of atom counts and molecule properties, with some subtle departures, most visible in $LogP$ and molecular weight.
We also compare 10,000 samples from a QM9-trained ‘Transdimensional Jump’ diffusion model (Campbell2023TransDimensionalGM), using a Kolmogorov-Smirnov based statistic ($1-KS_{D}$; see app. F.1) to quantify property agreement between samples and the data distribution (table 1). For all metrics, Branching Flows better matches the data distribution. We explore this further with visualizations in appendix fig 9.
4.2 Antibody Sequences
Table 2: Distribution agreement ($1-\mathrm{KS}_{D}$, higher is better) between properties of samples from each model vs. validation sequences. From left to right: Sequence Length, Diversity (within-samples min Cosine distance), Novelty (minimum edit distance to true sequence), CDR3 length, and AbNatiV ‘humanness’ score. Histograms of these properties are shown in Fig. 5.
Method Length Diversity Novelty CDR3 length AbNatiV
Oracle length 0.995 0.899 0.903 0.984 0.943
Branching Flows (ours) 0.986 0.894 0.888 0.989 0.931
Edit Flows 0.989 0.713 0.770 0.983 0.769
We investigate Branching Flows on a pure discrete sequence learning task: antibody amino acid sequence generation. Antibodies exhibit substantial length variation, most concentrated in the complementarity determining regions (CDRs) which are the primary determinants of antibody-antigen interaction. We compare a Branching Flows model to an ‘oracle-length’ discrete flow matching model, which is identical to the Branching Flows model except that it i) has no splits or deletions, and ii) is initialized with a ‘true’ length sampled from the data distribution. We also compare to a re-implementation of Edit Flows (Havasi2025EditFF). Full dataset details, Branching Flows specification, and model/training details are in Appendix F.2.
Figure 5 shows the distributions of statistics (Appendix F.2) from antibody sequences generated by both the Branching Flows and oracle-length models, perplexity-over-time training dynamics, and seqUMAP (hanke2022multivariate) representations of the 10 00010\,000 generated sequences. Distributional overlap with natural sequences is similar between Branching Flows and the oracle model, despite the oracle having additional information. The sequence-position-specific amino acid frequency distributions for all 20 amino acids in figure 10 shows that Branching Flows can learn strong position dependence despite not having any prior information about total length. Table 2 shows distribution matching statistics for the three models. In general, Branching Flows and the Oracle length model are very close across all metrics. Edit Flows performs well at length matching (both total and CDR3 length), but has lower distribution matching performance for metrics that consider the content of the generated sequences. App. fig. 11 explores these in more detail.
4.3 Protein Structure and Sequence
Figure 6: Branching Flows protein samples. Top left inset shows pseudoMSA scTM refolding scores, and selected samples are shown (rainbow colored for single chains, and per-chain for dimers). To the right or below each is shown an overlay of the sample and the refolded structure (grey). Specifically, the three worst scTMs are depicted (labelled a, b, and c on the inset and the main panel) showing that the low scTMs are driven by unpredictability due to flexible linkers, or incorrect dimerization prediction.
Starting with ChainStorm (OrestenChainStorm), a multimodal flow matching model that jointly generates protein backbones and amino acid sequence labels, as a base model, we finetuned two variants: For ‘BranchChain’, the designable residues are complete chains, randomly selected for masking during training, so it can generate chains unconditionally or it can generate chains conditioned on other fixed chains. BranchSegment’s conditioning mask during training is a random number of random-length contiguous segments, so multiple segments can be simultaneously designed during inference. Full setup and evaluation details are in Appendix F.3.
To investigate whether a Branching Flows protein model can generate samples with reasonable protein-like geometry, we sampled, without conditioning, 20 monomers and 20 dimers, discarding samples with chain breaks (one monomer and five dimers) and samples too short for meaningful evaluation (<25<25 residues in the longest chain: five monomers, and 1 dimer). Using the ‘pseudoMSA refolding’ strategy (OrestenChainStorm), we generated a pseudoMSA of 100 sequences for each design with ProteinMPNN (dauparas2022robust), and evaluate the ‘self-consistency TM scores’ of a single refold, via Boltz2 (passaro2025boltz) from this pseudoMSA.
Figure 6 shows the scTM scores, most very close to 1, showing that our designs have geometry consistent with Boltz2. We also depict selected examples, for monomers and dimers, of BranchChain generated proteins and their refolds, specifically showing the three cases with low scTMs. For the two poorly refolded monomers, this is because the generated structure had one or more flexible linkers, and the individual domains were well refolded. For the poorly refolded dimer, this was because BranchChain generated two disconnected chains, and Boltz2 placed these adjacent. Overall, BranchChain generates samples with a variety of protein chain lengths, with plausible backbone geometry. Further, as can be seen in many of the samples in figure 6, the propensity of the ChainStorm base model to sample homodimers is preserved. This is interesting given that ChainStorm’s single attention head that governs symmetric oligomer sampling is primarily driven by chain lengths which are fixed in ChainStorm, but now stochastically generated in BranchChain.
For illustration purposes, we also finetuned a BranchChain variant with only a single $x_{0}$ element per chain. Figure 7 shows steps along the sampling trajectory for a sample where two chains (blue and green) were fixed, and two chains were sampled. The orange chain, best seen through the lens of the model-predicted end state $\hat{x_{1}}$, highlights a key Branching Flows feature: at first, the designed chain converges upon two separate domains, with insufficient residues in between them for them to be linked. This is then progressively filled in by insertions, until a complete linker connects them as a single chain.
For BranchSegment we do not attempt quantitative evaluations, but we demonstrate that it is a promising solution to the ‘unknown length infix sampling’ problem. Taking a protein complex (PDB: 9IQP, shcheblyakov2025ultra) of a nanobody and the SARS-CoV-2 receptor binding domain (RBD), we mask only the ‘Complementarity Determining Region 3’ (CDR3) as designable, and we generate 20 samples from the RBD-bound complex, and 20 samples from the unbound nanobody alone. The sampled CDR3s had a broad length distribution (7 to 29 residues for bound, 7 to 40 unbound). For the bound samples, even though the side chain atom positions from the RBD were not provided to BranchSegment (it only sees backbone frames and amino acid labels) the designed CDR3s had no atomic clashes with the RBD side chain rotamers for 15 out of 20 of the generated CDR3s. Some of the unbound CDR3s appear structurally plausible, but are extremely unlikely to occur naturally, which is not surprising since neither the base model nor BranchSegment have any antibody or nanobody-specific training (see fig 8).
5 Related Work
5.1 Edit Flows
5.2 Transdimensional jumps
5.3 DrugFlow
6 Discussion
We have shown that Branching Flows is a capable ‘distribution-learner’ of multimodal sequences that vary in length, addressing a previously-unsolved fundamental problem in generative modeling. For example it is, to our knowledge, the first method that can sample an infix of unknown length from a multimodal sequence. For now, we restrict ourselves to demonstrating that Branching Flows works in a variety of domains, rather than extensive comparisons and benchmarks on downstream tasks, as performance on benchmarks is often more sensitive to practically useful but theoretically uninteresting domain-specific adjustments to the model and training details, which is orthogonal to the purpose of this manuscript.
For the QM9 small-molecule dataset, we compare samples from our Branching Flows-trained model with those from the Transdimensional Jump Diffusion model of Campbell2023TransDimensionalGM, showing favorable distribution matching for Branching Flows. We suspect this difference is attributable to two things. Firstly, transdimensional jump diffusion spawns elements at a position that is, in its conditional process, coupled to an element in the data distribution. While the conditional distribution of the spawn point is Gaussian conditioned on the corresponding $X_{1}$ element, the marginal distribution during inference is a (possibly infinite) mixture of Gaussians, which is challenging to approximate. Branching Flows avoids this by spawning new elements via a splitting mechanism that, conditioned on $X_{t}$ is independent of $X_{1}$. Secondly, transdimensional jump diffusion handles discrete atom types by converting them to continuous vectors. In other contexts, this does not compare favorably to discrete character diffusion models (sahoo2024simpleeffectivemaskeddiffusion).
For the antibody sequence modeling problem, we compare Branching Flows to an oracle-length model. During inference, the oracle length is sampled from the data distribution. Both the training dynamics of Branching Flows, and the ability to match the data distribution, are similar to a model that has been provided the true length. Edit Flows (Havasi2025EditFF), modeling variable length discrete sequences, is a natural comparator for the antibody sequence modeling problem. Here the model architectures are matched between Branching Flows and Edit Flows, where the only difference is that the Edit Flows insertion head is as large as the alphabet size, whereas Branching Flows only needs one splitting rate per element. While we see favorable performance of Branching Flows, one caveat is that this is our own Edit Flows implementation since, at the time of writing, there was no reference implementation provided by the authors.
Protein structure generation, as a field, has matured to the point where models are practically used in an experimental setting. It has recently become clear that in silico metrics, mostly relying on protein folding models, are poor predictors of experimental success (Mario2026Evaluating). As such, attempting to create and experimentally (not in silico) validate a state-of-the-art protein design model is beyond the scope of the current work. For now we content ourselves in showing that i) Branching Flows models generate samples that have geometry consistent with a state-of-the-art folding model, and ii) that they have new capabilities related to length variation that, if experimentally validated, would extend the space of problems approachable with protein design.
As far as we are aware ‘BranchSegment’ is the first generative model that can perform ‘unknown-length infix sampling’ for proteins, where the backbone is constrained either side of the designable region, and the length of the designed segment is determined by the model. In practice, current generative models must use a range of pre-sampled lengths, with post-sampling filters to select from these, and non-generative models could perform sequence search on a surrogate objective (e.g. via a folding model). Neither will provide distribution-matching, and both are computationally expensive. Further, the ‘pre-sampled length’ strategy required by fixed-length flow and diffusion models can combinatorially explode if there are multiple sequentially non-contiguous but spatially-interacting segments that need to be designed, which would be the case during protein-protein interface design or re-design.
One practically useful feature, demonstrated with ‘BranchChain’ and ‘BranchSegment’ is that Branching Flows can be finetuned into an existing model. Three days on a single GPU (NVIDIA RTX 6000 Ada) was sufficient. This suggests that, even though the base model was trained via fixed-length flow matching, the internal representations required to learn the branching and deletion process may already be present. This would be extremely useful if it extends to other models, allowing large models with high pretraining cost to be repurposed with Branching Flows.
If the elements are unordered, then Branching Flows can still be used, with adjustments, for example with the model being equivariant to permutations of the elements. However even for problems that are not naturally ordered (arguably the case for QM9 small molecules) initial exploration suggested that performance was much better when an order was imposed. We suspect this is due to model inadequacies when, for example, positional encoding is absent.
Branching Flows is, to our knowledge, the first ‘variable length Generator Matching’ method, and we anticipate that there is a large design space of alternatives, including methods that avoid the ‘splitting’ mechanism and instead spawn elements at points in space controlled by the model, or from some fixed base distribution.
Further work will explore the performance of Branching Flows in an array of different domains, and more systematically elucidate the design space of Branching Flows models and other variable length Generator Matching methods, comparing the benefits of different components of the process (e.g. branching-only vs deletion-only vs branching & deletion), as well as hyperparameters such as hazard distributions, anchor merging strategies, and the optimal choice of base process for different kinds of problems.
7 Acknowledgments
This project received support from the Swedish Research Council (2024-00390 and 2023-02516) and the Knut and Alice Wallenberg Foundation (2024.0039) to B.M. Development of key dependencies (e.g. https://github.com/MurrellGroup/Onion.jl) used in this work was enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. AI disclosure: GPT5 was used for initial drafting of some sections, which were rewritten and checked by the authors.
Appendix A $Z$ sampling schemes